The cost of two tables and three chairs is Ghc 705. If the table costs

Ghc 40 more than the chair, find the cost of the table and the chair.



1
Expert's answer
2022-05-16T16:36:49-0400

If the cost of the chair is "x" , so the cost of the table is "x+40" .

Because the cost of two tables and three chairs is GHC 705

we have: "2(x+40)+3x=705"

"2x+80+3x=705"

"5x=705-80"

"5x=625"

"x=125"

So, the cost of the chair is 125 and the cost of the table is "x+40=125+40=165"